# Let the Code LLM Edit Itself When You Edit the Code

### v1

@@ Abstract
In this work, we investigate a typical scenario in code generation where a developer edits existing code in real time and requests a code assistant, e.g., a large language model, to re-predict the next token or next line on the fly. Naively, the LLM needs to re-encode the entire KV cache to provide an accurate prediction. However, this process is computationally expensive, especially when the sequence length is long. Simply encoding the edited subsequence and integrating it to the original KV cache meets the temporal confusion problem, leading to significantly worse performance. We address this efficiency and accuracy trade-off by introducing Positional Integrity Encoding (PIE). Building upon the rotary positional encoding, PIE first removes the rotary matrices in the Key cache that introduce temporal confusion and then reapplies the correct rotary matrices. This process ensures that positional relationships between tokens are correct and requires only a single round of matrix multiplication. We validate the effectiveness of PIE through extensive experiments on the RepoBench-C-8k dataset, utilizing DeepSeek-Coder models with 1.3B, 6.7B, and 33B parameters. Our evaluation includes three real-world coding tasks: code insertion, code deletion, and multi-place code editing. Results demonstrate that PIE reduces computational overhead by over 85% compared to the standard full recomputation approach across all model sizes and tasks while well approximating the model performance.
@@ 3.1 Background
In this study, we aim to investigate a new scenario where the context $s$ is real-time edited by users, which makes it impossible for the KV cache to predict the correct next token without any modification. We can model such real-time context as a sequence of steps. Each step can be formulated as an action where tokens from position $i$ to $j$ in $s$ are edited, resulting in a modified sequence $s^{\text{edit}}=[w_{1},\dots,w_{i},a_{1},a_{2},\dots,a_{m},w_{j+1},\ldots,w_{n}]$, where $[a_{1},a_{2},\dots,a_{m}]$ represent the new inputs that replace $[w_{i},\dots,w_{j}]$. Our goal is to accurately and efficiently predict $w_{n+1}$ given by $\theta_{\text{LLM}}$ on $s^{\text{edit}}$. This problem is crucial in various scenarios. For instance, users can frequently edit their previous codes for different purposes and expect the code language model to swiftly adapt to these changes and predict the correct next line based on the updated information.
@@ 3.2 Positional Integrity Encoding (PIE)
Unlike the fully recomputation approach, the above calculation only requires a single round of matrix multiplication to directly modify the pre-computed KV cache, where the computational overhead can be considered negligible. By utilizing these transformations, we finally construct the edited KV cache as:
@@ 4 Experiments
We will answer each question with carefully designed experiments in the following sub-sections. Due to space limitations, more details are presented in Appendix A.
@@ 4.1 Experimental Setup
Our experiments are conducted on RepoBench-C-8k  (Liu et al., 2024a). This benchmark focuses on the prediction of the next line of code, given a set of in-file context (including import statements and preceding lines before the target line), and cross-file context (comprising snippets from other files parsed by import statements). The detailed statistics of RepoBench-C-8k is shown in Table 1. To effectively evaluate next-line prediction performance of code LLMs, we follow Liu et al. (2024a) to use three task settings: (1) Cross-File-First (XF-F): mask the first appearance of a cross-file line within a file; (2) Cross-File-Random (XF-R): mask a random and non-first occurrence of a cross-file line; (3) In-File (IF): mask an in-file line that does not involve any cross-file modules. Moreover, we carefully design three real-world scenarios covering code insertion, code deletion, and code edition to comprehensively examine our approach. See Appendix A for more detailed descriptions of tasks construction.
@@
In our experiments, we employ DeepSeek-Coder (Guo et al., 2024), a code LLM that achieves strong performance in handling repository-level code completion tasks. We use Transformers (Wolf et al., 2020) as our codebase. We benchmark our method on models of different sizes covering 1.3B, 6.7B, and 33B. During inference, the greedy decoding strategy is used to deterministically generate 64 tokens. For 1.3B and 6.7B models, all the experiments are conducted on a single NVIDIA A100 GPU. For 33B models, the time for encoding the context is conducted on two NVIDIA A100 GPUs and the full generation process is conducted on eight NVIDIA A100 GPUs. The first non-comment line in the output is truncated and used as the prediction. The batch size is set to 1. All experiments are repeated three times with different seeds and the averaged scores are reported.
@@ 4.2 Main Results
Table 2: Performance comparisons of insertion experiments. In this task, for each next-line prediction target, we insert several lines of code into its context randomly to simulate real-world scenarios. EM and ES denotes the Exact Match and Edit Similarity score respectively. All results demonstrate that our Positional Integrity Encoding approach brings substantial speed-ups without performance drops.
@@
33B Conflict Fast Encoding 0.38 2.51 120ms 0.59 2.4 122ms 0.68 2.09 122ms
@@
Table 4: Performance comparisons of edition experiments. In this task, for each next-line prediction target, we delete several lines of code of its context and simultaneously insert other lines of code randomly to simulate real-world scenarios. EM and ES denotes the Exact Match and Edit Similarity score respectively. All results demonstrate that our Positional Integrity Encoding approach brings substantial speed-ups without performance drops.
@@
1.3B Conflict Fast Encoding 8.8 50.08 23ms 13.83 54.01 22ms 11.29 51.99 22ms
@@
6.7B PIE 28.07 69.0 54ms 39.89 75.1 54ms 34.05 71.64 54ms
Python 33B Full-recomputation 35.75 73.46 2766ms 46.0 78.9 2759ms 39.75 75.12 2787ms
@@
1.3B Conflict Fast Encoding 5.87 31.46 22ms 8.01 32.74 23ms 10.2 34.34 23ms
@@
33B Conflict Fast Encoding 8.02 32.93 120ms 9.67 33.29 122ms 12.1 34.7 122ms
@@ 5 Conclusion
In this paper, we introduced Positional Integrity Encoding (PIE), a novel method designed to enhance the efficiency of large language models (LLMs) in the real-time editing setting. Our approach addresses the significant computational overhead associated with re-encoding contexts after small edits, a common scenario in interactive coding environments. Through extensive experiments, we demonstrated that PIE not only significantly reduces latency but also maintains high accuracy compared to the naive full re-computation method.

### v2

@@ Abstract
In this work, we investigate a typical scenario in code generation where a developer edits existing code in real time and requests a code assistant, e.g., a large language model, to re-predict the next token or next line on the fly. Naively, the LLM needs to re-encode the entire KV cache to provide an accurate prediction. However, this process is computationally expensive, especially when the sequence length is long. Simply encoding the edited subsequence and integrating it to the original KV cache meets the temporal confusion problem, leading to significantly worse performance. We address this efficiency and accuracy trade-off by introducing Positional Integrity Encoding (PIE). Building upon the rotary positional encoding, PIE first removes the rotary matrices in the Key cache that introduce temporal confusion and then reapplies the correct rotary matrices. This process ensures that positional relationships between tokens are correct and requires only a single round of matrix multiplication. We validate the effectiveness of PIE through extensive experiments on the RepoBench-C-8k dataset, utilizing DeepSeek-Coder models with 1.3B, 6.7B, and 33B parameters. Our evaluation includes three real-world coding tasks: code insertion, code deletion, and multi-place code editing. Results demonstrate that PIE reduces computational overhead by over 85% compared to the standard full recomputation approach across all model sizes and tasks while well approximating the model performance. Code is available at https://github.com/zhenyuhe00/PIE.
@@ 3.1 Background
In this study, we aim to investigate a new scenario where the context $s$ is real-time edited by users, which makes it impossible for the KV cache to predict the correct next token without any modification. We can model such real-time context as a sequence of steps. Each step can be formulated as an action where tokens from position $i$ to $j$ in $s$ are edited, resulting in a modified sequence $todo$ represent the new inputs that replace $[w_{i},\dots,w_{j}]$. Our goal is to accurately and efficiently predict $w_{n+1}$ given by $\theta_{\text{LLM}}$ on $s^{\text{edit}}$. This problem is crucial in various scenarios. For instance, users can frequently edit their previous codes for different purposes and expect the code language model to swiftly adapt to these changes and predict the correct next line based on the updated information.
@@ 3.2 Positional Integrity Encoding (PIE)
Unlike the full recomputation approach, the above calculation only requires a single round of matrix multiplication to directly modify the pre-computed KV cache, where the computational overhead can be considered negligible. By utilizing these transformations, we finally construct the edited KV cache as:
@@ 4 Experiments
We will answer each question with carefully designed experiments in the following sub-sections. We also conduct additional experiments in Appendix B.
@@ 4.1 Experimental Setup
Our experiments are conducted on RepoBench-C-8k  (Liu et al., 2024a). This benchmark focuses on the prediction of the next line of code, given a set of in-file context (including import statements and preceding lines before the target line), and cross-file context (comprising snippets from other files parsed by import statements). The detailed statistics of RepoBench-C-8k is shown in Table 1. To effectively evaluate next-line prediction performance of code LLMs, we follow Liu et al. (2024a) to use three task settings: (1) Cross-File-First (XF-F): mask the first appearance of a cross-file line within a file; (2) Cross-File-Random (XF-R): mask a random and non-first occurrence of a cross-file line; (3) In-File (IF): mask an in-file line that does not involve any cross-file modules. Moreover, we carefully design three real-world scenarios covering code insertion, code deletion, and code edition to comprehensively examine our approach. See Appendix A.1 for more detailed descriptions of tasks construction.
@@
In our experiments, we employ DeepSeek-Coder (Guo et al., 2024), a code LLM that achieves strong performance in handling repository-level code completion tasks (We also conduct experiments on CodeLlama (Roziere et al., 2023) in Appendix A.2). We use Transformers (Wolf et al., 2020) as our codebase. We benchmark our method on models of different sizes covering 1.3B, 6.7B, and 33B. During inference, the greedy decoding strategy is used to deterministically generate 64 tokens. For 1.3B and 6.7B models, all the experiments are conducted on a single NVIDIA A100 GPU. For 33B models, the time for encoding the context is conducted on two NVIDIA A100 GPUs and the full generation process is conducted on eight NVIDIA A100 GPUs. The first non-comment line in the output is truncated and used as the prediction. The batch size is set to 1. All experiments are repeated three times with different seeds and the averaged scores are reported.
@@ 4.2 Main Results
Table 2: Performance comparisons of insertion experiments. In this task, for each next-line prediction target, we insert several lines of code into its context randomly to simulate real-world scenarios. EM and ES denote the Exact Match and Edit Similarity score respectively. All results demonstrate that our Positional Integrity Encoding approach brings substantial speed-ups without performance drops.
@@
33B Conflict Fast Encoding 0.38 2.51 120ms 0.59 2.40 122ms 0.68 2.09 122ms
@@
Table 4: Performance comparisons of edition experiments. In this task, for each next-line prediction target, we delete several lines of code of its context and simultaneously insert other lines of code randomly to simulate real-world scenarios. EM and ES denote the Exact Match and Edit Similarity score respectively. All results demonstrate that our Positional Integrity Encoding approach brings substantial speed-ups without performance drops.
@@
1.3B Conflict Fast Encoding 8.80 50.08 23ms 13.83 54.01 22ms 11.29 51.99 22ms
@@
6.7B PIE 28.07 69.00 54ms 39.89 75.10 54ms 34.05 71.64 54ms
Python 33B Full-recomputation 35.75 73.46 2766ms 46.00 78.90 2759ms 39.75 75.12 2787ms
@@
1.3B Conflict Fast Encoding 5.87 31.46 22ms 8.01 32.74 23ms 10.20 34.34 23ms
@@
33B Conflict Fast Encoding 8.02 32.93 120ms 9.67 33.29 122ms 12.10 34.70 122ms
@@ 5 Conclusion
In this paper, we introduce Positional Integrity Encoding (PIE), a novel method designed to enhance the efficiency of large language models (LLMs) in the real-time editing setting. Our approach addresses the significant computational overhead associated with re-encoding contexts after small edits, a common scenario in interactive coding environments. Through extensive experiments, we demonstrated that PIE not only significantly reduces latency but also maintains high accuracy compared to the naive full re-computation method.
@@
Acknowledgements
We thank all the anonymous reviewers for the very careful and detailed reviews as well as the valuable suggestions. Their help has further enhanced our work. Di He is supported by National Science Foundation of China (NSFC62376007).
@@ A.1 Experimental Setup
A.2 Results on CodeLlama
Results of different code editing settings on CodeLlama (Roziere et al., 2023) are presented in Table 56 and 7 respectively. Similar to DeepSeek-Coder, CodeLlama with Positional Integrity Encoding demonstrates strong performance and fast speed across various editing scenarios. The metrics indicate that PIE effectively maintains prediction accuracy and addresses temporal confusion, ensuring the impact of minor edits is minimal.
Table 5: Performance comparisons of insertion experiments for CodeLlama. In this task, for each next-line prediction target, we insert several lines of code into its context randomly to simulate real-world scenarios. EM and ES denotes the Exact Match and Edit Similarity score respectively. All results demonstrate that our Positional Integrity Encoding approach brings substantial speed-ups without performance drops.
Model Method XF-F XF-R IF
EM ES Time EM ES Time EM ES Time
Python 7B Full-recomputation 25.72 66.49 561ms 38.87 73.81 564ms 33.55 70.26 562ms
7B Conflict Fast Encoding 14.49 55.10 34ms 19.32 57.34 34ms 15.49 53.05 34ms
7B PIE 25.42 66.40 50ms 38.76 73.83 50ms 33.30 70.11 50ms
Python 34B Full-recomputation 31.04 69.48 2013ms 42.80 76.27 2029ms 37.69 72.36 1999ms
34B Conflict Fast Encoding 11.56 50.23 113ms 16.06 52.59 110ms 11.58 48.36 112ms
34B PIE 30.53 68.79 123ms 42.73 76.40 119ms 37.51 72.25 123ms
Java 7B Full-recomputation 28.02 72.49 578ms 39.61 77.81 578ms 49.20 79.76 578ms
7B Conflict Fast Encoding 10.45 37.69 34ms 14.28 38.45 35ms 16.95 38.15 33ms
7B PIE 28.02 72.39 50ms 39.48 77.74 49ms 49.31 79.79 48ms
Java 34B Full-recomputation 32.16 75.09 2179ms 43.45 79.89 2184ms 52.43 80.97 2197ms
34B Conflict Fast Encoding 12.04 41.76 104ms 14.25 39.19 107ms 17.29 38.50 109ms
34B PIE 31.45 74.64 119ms 43.54 79.80 118ms 52.42 81.00 115ms
Table 6: Performance comparisons of deletion experiments for CodeLlama. In this task, for each next-line prediction target, we delete several lines of code of its context randomly to simulate real-world scenarios. EM and ES denote the Exact Match and Edit Similarity score respectively. All results demonstrate that our Positional Integrity Encoding approach brings substantial speed-ups without performance drops.
Model Method XF-F XF-R IF
EM ES Time EM ES Time EM ES Time
Python 7B Full-recomputation 25.72 66.49 561ms 38.87 73.81 564ms 33.55 70.26 562ms
7B Conflict Fast Encoding 10.14 52.80 30ms 15.24 57.88 31ms 13.04 54.53 30ms
7B PIE 25.52 66.47 43ms 38.88 73.87 42ms 33.51 70.18 42ms
Python 34B Full-recomputation 31.04 69.48 2013ms 42.80 76.27 2029ms 37.69 72.36 1999ms
34B Conflict Fast Encoding 12.15 54.76 88ms 16.40 59.29 94ms 13.78 55.90 93ms
34B PIE 31.01 69.44 100ms 42.76 76.13 100ms 37.60 72.38 102ms
Java 7B Full-recomputation 28.02 72.49 578ms 39.61 77.81 578ms 49.20 79.76 578ms
7B Conflict Fast Encoding 7.21 42.51 29ms 7.77 43.86 29ms 8.78 44.46 29ms
7B PIE 28.01 72.45 42ms 39.51 77.78 42ms 49.30 79.80 42ms
Java 34B Full-recomputation 32.16 75.09 2179ms 43.45 79.89 2184ms 52.43 80.97 2197ms
34B Conflict Fast Encoding 8.76 45.44 92ms 8.67 45.57 92ms 10.11 46.64 95ms
34B PIE 32.12 75.02 99ms 43.33 79.86 98ms 52.36 80.96 107ms
Table 7: Performance comparisons of edition experiments for CodeLlama. In this task, for each next-line prediction target, we delete several lines of code of its context and simultaneously insert other lines of code randomly to simulate real-world scenarios. EM and ES denote the Exact Match and Edit Similarity score respectively. All results demonstrate that our Positional Integrity Encoding approach brings substantial speed-ups without performance drops.
Model Method XF-F XF-R IF
EM ES Time EM ES Time EM ES Time
Python 7B Full-recomputation 25.72 66.49 705ms 38.87 73.81 706ms 33.55 70.26 713ms
7B Conflict Fast Encoding 19.97 61.46 34ms 29.79 67.31 34ms 24.70 63.28 34ms
7B PIE 24.97 66.04 54ms 38.21 73.40 54ms 32.84 69.66 54ms
Python 34B Full-recomputation 31.04 69.48 2574ms 42.80 76.27 2569ms 37.69 72.36 2581ms
34B Conflict Fast Encoding 21.91 61.78 91ms 30.87 67.16 91ms 25.76 63.15 90ms
34B PIE 29.92 68.7 126ms 41.74 75.37 126ms 37.01 71.76 127ms
Java 7B Full-recomputation 28.02 72.49 733ms 39.61 77.81 736ms 49.20 79.76 728ms
7B Conflict Fast Encoding 18.78 57.38 34ms 25.79 60.97 35ms 32.60 62.66 33ms
7B PIE 27.43 71.60 52ms 38.89 76.96 53ms 48.60 78.94 53ms
Java 34B Full-recomputation 32.16 75.09 2726ms 43.45 79.89 2807ms 52.43 80.97 2727ms
34B Conflict Fast Encoding 21.25 59.70 104ms 27.83 61.91 107ms 32.98 62.03 109ms
34B PIE 30.82 73.84 126ms 42.71 78.94 119ms 51.64 80.15 125ms
Appendix B Additional Experiments
B.1 Code Generation Tasks
To further validate the effectiveness of PIE, we conduct experiments on code generation tasks, using HumanEval Chen et al. (2021) and its C++ version from HumanEval-X Zheng et al. (2023). We randomly choose three consecutive lines in the prompt as the user inserted code and report Pass@1 for different methods. It’s important to note that the context (prompt) in HumanEval consists of only a few lines (with an average of 15 lines), making each edit semantically significant. The results in Table 8 demonstrate that PIE significantly outperforms the baseline methods, approaching the performance of full recomputation. Note that we also include a reuse baseline where the pre-edit context is reused without modification. Given the importance of these edits in such short contexts, updating the KV cache with PIE becomes essential for maintaining high prediction accuracy, whereas the performance of the reuse baseline drops substantially.
B.2 Contextually Related Edits
We design a contextually related insertion setting, where we use Edit Distance to identify lines in the context that are similar to the target line, simulating user insertions. We also increase the number of edited lines. We conduct experiments on the Cross-File-First setting of RepoBench-C-8k, keeping all other settings the same as Section 4.The results in Table 9 demonstrate that PIE is robust in handling contextually related edits and outperforms the baseline methods, approaching the performance of full recomputation. Note that we also include a reuse baseline where the pre-edit context is reused without modification.
B.3 Multi-Place Edits
Users may edit a function definition followed by updates to all associated call sites. To simulate this scenario, we treat each line in the context as a potential site and use Edit Distance to identify multiple sites in the context that are similar to the target line, thereby simulating multi-place insertions. We conduct experiments in this contextually related multi-place editing setting on the Cross-File-First task of RepoBench-C-8k. The results in Table 10 demonstrate that PIE is robust in handling multi-place edits and outperforms the baseline methods.
Table 8: Performance of DeepSeek-Coder 6.7B on HumanEval and HumanEval-X
Method Pass@1
Python C++
Full-recomputation 49.4 50.3
Conflict Fast Encoding 11.0 15.5
Reuse baseline 17.7 29.2
PIE 43.9 49.7
Table 9: Performance of DeepSeek-Coder 6.7B on contextually-related insertion experiments
Model Method 5 lines 7 lines 11 lines 13 lines 15 lines
EM ES EM ES EM ES EM ES EM ES
Python 6.7B Full-recomputation 28.95 70.11 28.95 70.11 28.95 70.11 28.95 70.11 28.95 70.11
6.7B Conflict Fast Encoding 7.92 41.56 6.39 38.56 4.78 34.91 4.11 33.34 3.43 32.15
6.7B Reuse baseline 25.58 68.23 25.29 68.03 25.23 67.95 25.13 67.83 24.71 67.74
6.7B PIE 27.40 69.10 27.23 68.94 27.04 68.80 26.92 68.8 26.86 68.63
Java 6.7B Full-recomputation 32.51 75.56 32.51 75.56 32.51 75.56 32.51 75.56 32.51 75.56
6.7B Conflict Fast Encoding 2.81 15.46 1.89 12.16 1.10 7.89 0.87 6.52 0.74 5.57
6.7B Reuse baseline 25.99 73.29 25.72 73.17 25.18 73.03 25.15 72.99 24.99 72.84
6.7B PIE 30.04 74.46 30.00 74.44 30.13 74.47 29.87 74.38 29.76 74.29
Table 10: Performance of DeepSeek-Coder 6.7B on contextually-related multi-place insertion experiments
Model Method 5 lines 7 lines 11 lines 13 lines 15 lines
EM ES EM ES EM ES EM ES EM ES
Python 6.7B Full-recomputation 28.95 70.11 28.95 70.11 28.95 70.11 28.95 70.11 28.95 70.11
6.7B Conflict Fast Encoding 7.43 39.95 5.83 37.00 4.06 33.02 3.48 31.71 2.97 30.59
6.7B Reuse baseline 24.61 67.43 24.44 67.17 23.84 66.77 23.62 66.63 23.49 66.55
6.7B PIE 26.97 68.85 26.77 68.67 26.33 68.34 26.15 68.222 25.9 68.23
Java 6.7B Full-recomputation 32.51 75.56 32.51 75.56 32.51 75.56 32.51 75.56 32.51 75.56
6.7B Conflict Fast Encoding 1.69 10.91 1.06 7.66 0.43 4.26 0.28 3.36 0.22 2.81
6.7B Reuse baseline 24.46 72.35 23.99 72.05 23.11 71.69 22.97 71.56 22.75 71.38
6.7B PIE 29.56 74.14 29.26 73.91 29.03 73.81 28.83 73.68 28.56 73.59